# Scalable Cosmic AI Inference using Cloud Serverless Computing with FMI

### v2

@@ 5.2 Implementation
We used Python 3.11 with PyTorch 2 as the core framework. Also, NumPy 1.2 and Pandas 2.2 for data analytics. Additionally, Timm 0.4.12 was independently installed to provide access to specific model architectures compatible with PyTorch Paszke et al. (2019). The FMI and Boto3 libraries were used on AWS for the CAI implementation.
@@ 5.4 Evaluation Metrics
For CAI, since it is running parallel jobs, we calculate the inference time as the time needed to finish the Distributed Model Inference state (Figure 3) from the AWS step machine. The throughput is also calculated similarly, based on total bits calculated by all jobs divided by the total time to finish this state.
@@ 5.5 Experiment Setup
We compare the following processors: a personal laptop, HPC CPU, HPC GPU, and our proposed CAI. Figure 5 shows their specifications and Table 3 lists the maximum dataset size they can handle. A batch size of 512 is used throughout the experiment. Each experiment combination is repeated three times and the average metrics are reported.
Table 3: Processors’ maximum dataset size.
Processor Maximum Dataset Size (GB) Number of Images
Personal Laptop 8 418,323
HPC GPU 12.6 (Full dataset) 659,857
HPC CPU 12.6 (Full dataset) 659,857
CAI 12.6 (Full dataset) 659,857
First, we investigate how varying dataset sizes affect inference time, benchmarking the scalability of each device. The complete dataset is partitioned into incremental sizes: 1 GB, 2 GB, 4 GB, 6 GB, 8 GB, 10 GB, and 12.6 GB (the full dataset). This allows us to evaluate the total inference time on different devices and identify the maximum dataset size each device can handle. Second, we examine the impact of batch size on throughput (bps) across different devices. Batch sizes of 32, 64, 128, 256, and 512 are tested using a fixed dataset size of 1 GB. Finally, we evaluate the choice of partition size for CAI, using 25MB, 50MB, 75MB, and 100MB. This shows how different partition size choices could impact the inference time and throughput.
@@ 5.6 Performance Results and Analysis
Figure 6 shows that the inference time increases with data size. The personal laptop could not perform inference on the larger dataset sizes and is much slower due to memory limitations. The HPC with GPU is way faster than with CPU and personal laptop. However, CAI maintains a considerably stable inference time and performs best, due to its parallel executions, which is more evident with larger data sizes.
Model inference on the total data takes 1793 s, 140.8 s, and 28 s on the HPC CPU, HPC GPU, and CAI, respectively. The personal laptop could run to a maximum of 8GB and took 6283 s on average. This showcases personal devices are limited by memory and not feasible for the long run-times with astronomy data. Our proposed CAI is an attractive approach due to the faster inference speed despite running on CPU only.
Figure 7 shows the throughput of the devices for 1GB data size. 1GB was chosen due to the limitation of the personal laptop taking significant time with small batch size and larger data. Both the HPC GPU and CAI have significantly higher throughputs. CAI has the highest throughput (bps) for batch sizes 32, 64, and 128 (0.25B bps, 0.34B bps, and 0.42B bps respectively). The HPC GPU has the highest throughput (bps) for batch sizes 256 and 512 (0.42B bps and 0.61B bps). After CAI and the HPC GPU, the HPC CPU tends to have the next highest throughput, and then lastly, the personal laptop.
@@ 5.7 Serverless Computing for Scalable Cosmic AI
Figure 9: CAI Average Throughput for Different Data Sizes. These results are over concurrent jobs and use batch size 512) for different partition sizes.
@@
Runtime Cost analysis: AWS Lambda costs are based on requests and the duration time related to the execution of the Lambda function. For CAI, we use three AWS Lambda functions: 1) data-parallel-init-fmi, 2) cosmic-executor, and 3) resultSummary. The material cost is the number of executions of the cosmic executor. The costs for 1) and 2) are marginal based on the fact that these are only single executions for a given job (i.e., 25GM, 50GB, 75GM, or 100GB partition(s)), and each of these lambda functions is provisioned very conservatively with 128MB memory and 512MB ephemeral storage. AWS published guidance indicates Lambda costs for the CAI use case to be based on architecture (x86 or ARM), memory, ephemeral storage, and data transfer for CAI. We have provisioned our cosmic-executor AWS Lambda function as follows:
@@
The cost of invoking the CAI cosmic-executor AWS Lambda function is \$0.00001667 per GB-second of computation time. It should also be mentioned that we have small S3 costs related to storage (our S3 Bucket is around 50GB) and data transfer. For the experiments conducted in the aggregate, this amounts to \$0.023 per GB, and data transfer costs \$0.005 per 1,000 requests. This cost is non-material, similar to costs related to executing the data-parallel-init-fmi and resultSummary AWS Lambda function. Figure 3 shows that our framework calls the Lambda function during initialization, parallel processing, and summarization. Table 5 summarizes some example cases to estimate the computation cost for our task. The number of requests is how often the Lambda function was called, which is the number of concurrent jobs (data divided by partition size).
@@
Table 5: An estimation of AWS computation cost for inference on the total dataset. Cost in US cents is $requests\times duration(s)\times memory(GB)$.
All this amounts to under five US dollars to execute the experiments. As mentioned, CAI uses the FMI library for summation and the all-reduce collective to generate metrics related to the average time required to execute the Lambda function pairs. Collective operations are used as a portable interface for parallel programming. Replacing send/receive operations with collectives increases performance and facilitates debugging and maintenance. Collectives also lead to a ”…”division of labor: the programmer thinks in terms of these primitives and the library is responsible for implementing them efficiently.”Peter Sanders (2009). For example, we determined the average summation time to be measured across twelve AWS Lambda functions for a 1 GB partition processing inference to be 44.012 seconds without resorting to manual calculation or another process to calculate this result. This accurately measures the processing time across AWS Lambda functions and allows researchers a clear view of execution time as a function of cost. One novelty aspect of CAI is related to the costs of executing astronomy inference compared to similar costs on HPC clusters or other ”rack and stack infrastructure.” Coupling this with the integration of FMI allows researchers to utilize a configured lambda function’s processing capability fully. With FMI, we can configure a fleet of AWS Lambda functions and bypass a subset of the costs related to invocation by using communication between AWS Lambda functions in the form of Nat Traversal or point-to-point communication.
@@ 6 Conclusion and Future Work
To showcase the attractive capabilities of CAI, we used a recent foundation model, fine-tuned for redshift prediction, in our experiments. Comprehensive evaluations across varying dataset sizes and computing devices demonstrate CAI’s robustness in scaling almost linearly on the inference and high throughput for large-scale astronomical image datasets. Note that this framework can be applied to additional critical astronomy inference tasks, such as morphology classification and star formation history. Future work will also involve testing other foundation models developed for astronomical images and spectra. Moreover, we plan to integrate the FMI communication interface into the open-source Cylon Perera et al. (2023) library. This will build upon the high-performance communication libraries currently supported (e.g., MPI, UCC/UCX, Gloo).
@@ 7 Acknowledgement
This work is partly supported by NSF grant CCF-1918626 Expeditions: Collaborative Research: Global Pervasive Computational Epidemiology, and NSF Grant 2200409 for CyberTraining:CIC: CyberTraining for Students and Technologies from Generation Z.

### v3

@@ 5.2 Implementation
We used Python 3.11 with PyTorch 2 as the core framework. Also, NumPy 1.2 and Pandas 2.2 for data analytics. Additionally, Timm 0.4.12 was independently installed to provide access to specific model architectures compatible with PyTorch (Paszke et al., 2019). The FMI and Boto3 libraries were used on AWS for the CAI implementation.
@@ 5.4 Evaluation Metrics
For CAI, since it is running parallel jobs, we calculate the inference time as the time needed to finish the Distributed Model Inference state (Figure 3) from the AWS step machine. The throughput is also calculated similarly, based on the total bits calculated by all jobs divided by the total time to finish this state.
@@ 5.5 Experiment Setup
We compare the following processors: a personal laptop, HPC CPU, HPC GPU, and our proposed CAI. Figure 5 shows their specifications, and Table 3 lists the maximum dataset size they can handle. A batch size of 512 is used throughout the experiment. Each experiment combination is repeated three times, and the average metrics are reported.
Device Maximum Data (GB) #Cores RAM (GB) #Images
Laptop 8 16 16 418K
HPC GPU 12.6 128 400 660K
HPC CPU 12.6 128 400 660K
CAI 12.6 6 10 660K
Table 3: Hardware details and maximum data processed for each device. The original full dataset is 12.6 GB.
First, we investigate how varying dataset sizes affect inference time, benchmarking the scalability of each device. The complete dataset is partitioned into incremental sizes: 1 GB, 2 GB, 4 GB, 6 GB, 8 GB, 10 GB, and 12.6 GB (the full dataset). This enables us to evaluate the total inference time on various devices and determine the maximum dataset size each device can handle. Second, we examine the impact of batch size on throughput (bps) across different devices. Batch sizes of 32, 64, 128, 256, and 512 are tested using a fixed dataset size of 1 GB. Finally, we evaluate the choice of partition size for CAI, using 25MB, 50MB, 75MB, and 100MB. This shows how different partition size choices could impact the inference time and throughput.
@@ 5.6 Performance Results and Analysis
Figure 6 shows that the inference time increases with data size. The personal laptop could not perform inference on the larger dataset sizes and is much slower due to memory limitations. The HPC with GPU is way faster than with a CPU and a personal laptop. However, CAI maintains a considerably stable inference time and performs best, due to its parallel executions, which is more evident with larger data sizes.
Model inference on the total data takes 1793 s, 140.8 s, and 28 s on the HPC CPU, HPC GPU, and CAI, respectively. The personal laptop could run with a maximum of 8GB and took an average of 6283 seconds. This showcases that personal devices are limited by memory and not feasible for the long run-times with astronomy data. Our proposed CAI is an attractive approach due to the faster inference speed despite running on a CPU only.
Figure 7 shows the throughput of the devices for 1GB data size. 1GB was chosen due to the limitation of the personal laptop, taking significant time with a small batch size and larger data. Both the HPC GPU and CAI have significantly higher throughputs. CAI has the highest throughput (bps) for batch sizes 32, 64, and 128 (0.25B bps, 0.34B bps, and 0.42B bps, respectively). The HPC GPU has the highest throughput (bps) for batch sizes 256 and 512 (0.42B bps and 0.61B bps). After the CAI and the HPC GPU, the HPC CPU tends to have the next highest throughput, followed by the personal laptop.
@@ 5.7 Serverless Computing for Scalable Cosmic AI
Figure 9: CAI Average Throughput for Different Data Sizes. These results are over concurrent jobs and use a batch size of 512 for different partition sizes.
@@
With AWS Lambda, it is quite trivial to scale out horizontally to handle large dataset sizes. Figure 10 details the number of invocations and inference duration to illustrate scalability. The number of invocations is determined by dividing the partition size by the dataset size.
Figure 10: CAI: number of AWS Lambda Function invocations for different data sizes. Similar to figure 9, results use a batch size of 512 and are displayed across partition sizes.
Runtime Cost analysis: AWS Lambda costs are based on Table 4. For CAI, we use three AWS Lambda functions: 1) data-parallel-init-fmi, 2) cosmic-executor, and 3) resultSummary. The material cost is the number of executions of the cosmic executor. The costs for 1) and 2) are marginal based on the fact that these are only single executions for a given job (i.e., 25 GB, 50 GB, 75 GB, or 100 GB partition(s) and larger scaling tests for the 100GB partition for 256 GB, 512 GB, 768GB and one terabyte data sizes), and each of these lambda functions is provisioned very conservatively with 128MB memory and 512MB ephemeral storage. AWS published guidance indicates Lambda costs for the CAI use case to be based on architecture (x86 or ARM), memory, ephemeral storage, and data transfer for CAI. We have provisioned our cosmic-executor AWS Lambda function as detailed in Table 5.
Detail Cost
x86 Arch: First 6 Billion GB-seconds/month;
Duration: \$0.0000166667 per GB-second
\$0.20 per 1M requests
Memory (MB): 10,240 \$0.0000000021 per 1ms
Ephemeral Storage \$0.0000000309 for every GB-second
Table 4: AWS Lambda Costs for CAI Executions
@@
The cost of invoking the CAI cosmic-executor AWS Lambda function is \$0.00001667 per GB-second of computation time. It should also be mentioned that we have small S3 costs related to storage (our S3 Bucket is around 50GB) and data transfer. For the experiments conducted in the aggregate, this amounts to \$0.023 per GB, and data transfer costs \$0.005 per 1,000 requests. This cost is non-material, similar to costs related to executing the data-parallel-init-fmi and resultSummary AWS Lambda function. Figure 3 shows that our framework calls the Lambda function during initialization, parallel processing, and summarization. Table 6 summarizes some example cases to estimate the computation cost for our task. The number of requests represents the frequency at which the Lambda function was called, which is equivalent to the number of concurrent jobs (data divided by partition size).
@@
100MB 10240 23.0 s 7.0GB 27.48
Table 6: An estimation of AWS computation cost for inference on the total dataset up to one terabyte data size. Cost in US cents is $requests\times duration(s)\times memory(GB)\times 0.00001667$.
All this amounts to under five US dollars to execute the experiments for all but the largest experiment. One novelty aspect of CAI is related to the costs of executing astronomy inference compared to similar costs on HPC clusters or other ”rack and stack infrastructure.”
5.8 Large Scale Experiment and Analysis
We demonstrate the scalability of our CAI framework by conducting experiments on significantly larger datasets, up to 1 Terabyte (1024 GB). This is achieved by repeatedly sampling from the original 12.6 GB data partitions. We then compare CAI’s performance with that of a traditional High-Performance Computing (HPC) environment. The experiments were performed on HPC CPUs to make a fair comparison with CAI, which is also CPU-only. Both setup uses the 100 MB file partition for consistency. Table 7 shows the function invocations for each data size. This is determined by the number of partitions we need to process to reach the target data size. Calculated by dividing the data size in MB by the partition size, 100 MB.
5.8.1 HPC CPU Setup:
The HPC environment uses the SLURM job scheduler (Jette and Wickberg, 2023) to submit a large number of concurrent jobs. Similar to CAI, each job processes a single file partition on a single CPU core. The ‘srun‘ Python command then executes the inference on all allocated CPU cores. The concurrency limit for CPU cores on the HPC was 6000; beyond that, it’ll wait for the running jobs to finish. The virtual environment uses the same Python library versions as the AWS CAI framework. Each HPC node had maximum 96 cores and 750 GB of memory.
Data (GB) 100 256 512 768 1024
#Parallel Functions 103 2622 5243 7865 10486
Table 7: Setup summary of the extended scaling analysis: run up to 1TB data over 10486 parallel CAI Lambda functions on AWS and 10486 parallel functions on the HPC system.
5.8.2 Results and Discussion:
Figure 11 shows the average inference time for both the CAI framework and the HPC CPU environment across various large datasets, from 100 GB to 1024 GB. The results clearly demonstrate the superior scalability of the CAI framework.
Figure 11: Extended large-scale scaling experiment: Comparing CAI Lambda and the HPC system on inference time with up to 1 TB of data and 10,486 parallel functions.
CAI maintains a consistently low inference duration across the entire range of dataset sizes. This linear scalability is a direct result of its serverless architecture, which allows a high degree of parallelization by processing multiple data partitions concurrently. The inference time is primarily dependent on the time it takes to process a single data partition, not the total dataset size.
@@ 6 Conclusion and Future Work
To showcase the attractive capabilities of CAI, we used a recent foundation model, fine-tuned for redshift prediction, in our experiments. Comprehensive evaluations across varying dataset sizes and computing devices demonstrate CAI’s robustness in scaling almost linearly on the inference and high throughput for large-scale astronomical image datasets. Note that this framework can be applied to additional critical astronomy inference tasks, such as morphology classification and star formation history. Future work will also involve testing other foundation models developed for astronomical images and spectra.
In the future, we plan to introduce integration with FMI, FaaS Message Interface (Böhringer, 2022) for high-performance communication between AWS Lambda functions and to introduce support for collective operations. Direct communication between AWS Lambda functions over TCP sockets is not supported, based on published capabilities and the underlying design of Lambda functions by Amazon, which are intended to process lightweight event notifications or request payloads. To elaborate, AWS Lambda communication is capped at 6 MB for synchronous request/response, 20 MB for streamed data, and 256 KB for asynchronous. Direct communication offers performance benefits related to message passing between Lambda functions compared to JSON payloads and other solutions that introduce intermediary data storage to overcome communication performance and size limitations. With the FMI integration, functions will be able to participate in point-to-point, high-performance message passing and perform collective operations during later stages (e.g., model inference). For our scaling experiments with data sizes of 256, 512, 768, and one terabyte, it was necessary to use intermediate storage based on inbound and outbound payload size limitations. FMI Böhringer et al. (2023) demonstrates significant improvements ranging from 105 to 1025 times compared to storage mechanisms like AWS S3, Redis, or AWS DynamoDB, we can enhance both costs and performance. Furthermore, we can tackle the crucial challenge of communication performance in AI/ML inference on AWS Lambda.
@@ 7 Acknowledgement
This work is partly supported by NSF-Simons AI Institute for Cosmic Origins (CosmicAI: Grant 2421782) Seed Grant. NSF CF-1918626 Expeditions: Collaborative Research: Global Pervasive Computational Epidemiology, and NSF Grant 2200409 for CyberTraining:CIC: CyberTraining for Students and Technologies from Generation Z.